### v1

# Scattering of CO from vacant-MoSe2 with O adsorbates: is CO2 formed?

Abstract
Using ab initio molecular dynamics (AIMD) simulations, based on density functional theory that also accounts for van der Waals interactions, we study the oxidation of gas phase CO on MoSe2 with a Se vacancy and oxygen coverage of 0.125 ML. In the equilibrium configuration, one of the O atoms is adsorbed on the vacancy and the other one atop one Se atom. Recombination of the CO molecule with the second of these O atoms to form CO2 is a highly exothermic reaction, with an energy gain of around 3 eV. The likeliness of the CO oxidation reaction on this surface is next examined by calculating hundreds of AIMD trajectories for incidence energies that suffice to overcome the energy barriers in the entrance channel of the CO oxidative recombination. In spite of it, no CO2 formation event is obtained. In most of the calculated trajectories the incoming CO molecule is directly reflected and in some cases, mainly at low energies, the molecules remain trapped at the surface but without reacting. As important conclusion, our AIMD simulations show that the recombination of CO molecules with adsorbed O atoms is a very unlikely reaction in this system, despite its large exothermicity.
[…]
![[Uncaptioned image]](extracted/5871042/figures/TOC-1.png)
1 Introduction
In the last years, the possibility of using two-dimensional (2D) materials for catalysis 1, 2, 3, 4, 5, 6 and gas sensing 7, 8, 9, 10 has been broadly explored. Different families of 2D materials, like transition metal carbides or nitrides 11 and transition metal dichalcogenides (TMD) 12, 13, 14, 15, 16, 17, 18, have been considered for these purposes. The motivation roots in the exceptional properties that these materials exhibit for such applications, as for example, a large surface to volume ratio and good electronic and optical properties. Importantly, their properties can be tuned by several procedures, including application of strain 19, application of an external electric field 20, creation of vacancies 21, 22, 23, 24, 25, stacking different two-dimensional materials 26, 10, and even by considering non 2D configurations 27, 28, 29. The latter permits to engineer new 2D materials with the desired properties. For gas sensing and electrocatalysis applications, TMD semiconductors, in which we focus on in this work, offer the advantage of being easily implemented with electrodes16, while overcoming the zero band-gap problem of graphene that complicates its implementation in electronic circuits 30.
Focusing on the usage of TMDs as catalysts for diverse chemical processes, their applicability is promising. There are various studies exploring the catalytic properties of TMDs for the hydrogen evolution reaction (HER) 31, 32, 33, 24, 3, 34, 25 and the oxygen evolution reaction (OER) 35, 36, 37. These are the two reactions involved in water splitting, a key process for sustainable energy production. In the case of HER, TMD materials appear as good alternatives to precious metals such as platinum. However, pristine TMDs exhibit poor performance for OER due to the large binding energies of intermediate products on the TMD surface 35. Nonetheless, both theory and experiments reveal that the efficiency of these two reactions can be enhanced at the edges of the nanosheets 37, 36, 3, around chalcogen and metal vacancies 31, 32, 24, and by doping the nanosheet 33. Other chemical reactions that are the focus on theoretical studies of the TMD catalytic properties include, for example, CO2 reduction to produce acids and other valuable products 38, 39, 40, 41, 42, 43, 1, 2, 6, N2 reduction to produce ammonia 44, production of ethanol from syngas 45, and also CO oxidation on MoS2 doped with metal atoms 46, 47, on PtTe2 in the presence of O2 48 and on the Janus WSSe surface 49.
In the aforereviewed bibliography, the catalytic properties are commonly determined in terms of the adsorption energies and minimum energy paths of the reaction considered. Being the latter the common procedure, the purpose and novelty of our work are to explore the catalytic properties of MoSe2 for CO2 formation but from a different perspective. In particular, we focus on the dynamics of the recombination process between one O atom adsorbed on the MoSe2 surface with a Se vacancy and a gas-phase CO molecule. This kind of reaction process, which can occur upon a single collision of the gas species with either the adsorbate (direct Eley-Rideal) or the surface (indirect Eley-Rideal) or after few collisions that serve to slow down the incoming atom/molecule (hot atom/molecule abstraction), are expected to efficiently compete with the thermal Langmuir-Hinshelwood recombination when the adsorbate is strongly bound to the surface. This is precisely the case in the dissociative adsorption of O2 on the MoSe2 surface with a Se vacancy. DFT calculations with the revPBE-vdW exchange-correlation functional developed by Dion et al. 50 showed that the dissociative adsorption of O2 is exothermic, being the adsorption energies of the nascent O atoms that adsorb at different sites eV and eV 51. Our new calculations for the recombinative formation of CO2 via an Eley-Rideal mechanism show that the reaction is energetically favored and thus, motivated us to study the process with ab initio molecular dynamics simulations (AIMD). Interestingly, none of the incoming CO molecules is able to abstract an O adatom and form CO2, even if the molecule has enough energy to overcome the entrance barriers for the reaction. These results remark the importance of going beyond static calculations of the reaction energetics in determining the catalyzing character of the material under consideration.
The paper is organized as follows, the general computational settings employed to both model the system and simulate the reaction dynamics with the AIMD methodology are detailed in the Theoretical Methods section. The Results and Discussion section starts with the analysis of the energetics of the main reactions that may occur in the system for the incidence conditions considered, i.e., a gas-phase CO impinging on the MoSe2 surface with a Se vacancy and an oxygen coverage of 0.125 ML. Next, the AIMD results are presented and discussed in the light of the information extracted from the in depth analysis of the trajectories. The summary and conclusions of our work are given in the last section.
2 Theoretical Methods
2.1 General DFT computational settings
All DFT calculations are performed with the Vienna ab initio simulation package (vasp) 52 using the revPBE-vdW exchange-correlation functional proposed by Dion et al. 50. The ionic cores are described with the projector augmented wave (PAW) method 53 implemented in VASP 54. Specifically, we use the PAW potentials for Mo, Se, O, and C that have 14, 6, 6, and 4 valence electrons, respectively. The Kohn-Sham states are expanded in a plane-wave basis set using an energy cutoff of 700 eV. The Brillouin zone integration is performed with a -centered 441 Monkhorst–Pack grid of special -points 55. A Gaussian smearing of 0.1 eV is used for electronic state occupancies.
Calculations are performed for the phase of the MoSe2 monolayer with a small concentration of non-interacting Se vacancies (denoted as MoSe2-vSe hereafter). Following ref 51, the latter is adequately described by a 44 surface supercell that contains a single Se vacancy and is separated from its periodic image by 23 Å of vacuum along the surface normal. The value of the hexagonal lattice constant is =3.386 Å 51. The O2 molecule is dissociatively adsorbed in the configuration analyzed in ref 51. Namely, one of the O atoms is adsorbed on top of the Se vacancy and the other one on top of the closest Se atom. The O atom in the vacancy is located at Å below the surface, the O atom on top of the Se atom is located at =1.57 Å above the surface, and the distance between them is d(O-O)=4.01 Å. Dissociation of the O2 molecule in this configuration is exothermic by 3.3 eV. We denote this structure as 2O-MoSe2-vSe.
2.2 AIMD Initial Conditions
The interaction dynamics of CO with the 2O-MoSe2-vSe surface is studied with total energy-conserving AIMD simulations for an initial substrate temperature of 100 K. The initial positions and velocities of the atoms in the substrate are randomly taken from a canonical ensemble of configurations. The latter is generated by thermalizing the 2O-MoSe2-vSe substrate at 100 K during more than 10 ps using the Nosé-Hoover thermostat 56 implemented in VASP. The use of constant-energy AIMD simulations with an appropriate set of initial positions and velocities constitutes a reasonable approximation to incorporate surface temperature effects during the short simulation time of interest (up to 1.0–1.4 ps in our case) 57. This is a common procedure in gas-surface dynamics simulations 58, 59, 60, 61, 57, 62, 63, 64, 65, 66, 67.
The CO molecule starts with its center of mass (CM) located at a distance Å from the 2O-MoSe2-vSe surface with the vibrational zero-point energy and zero rotational energy. The initial molecular orientation and CM position in the -plane parallel to the surface are randomly sampled. The (,)-sampling however focuses on the simulation cell area that contains the O adatoms with the purpose of optimizing the computational effort (see below). The molecule impinges perpendicular to the surface (i.e., normal incidence). Simulations are performed for five different initial translational energies, namely, =0.2, 0.4, 0.6, 0.8, and 1.0 eV. For each , 200 AIMD trajectories are run using the Beeman predictor-corrector algorithm with an integration time step of 1 fs. At the end of each AIMD trajectory, the impinging CO is considered reflected when its center of mass reaches a distance from the Se topmost layer Å with positive velocity along the surface normal (). It is considered trapped if after reaching the maximum integration time of 1 ps, the previous condition was not met. In the case of recombinative CO2 formation, the same criteria is used to distinguish between CO2 desorption and trapping.
3 Results and Discussion
The calculated reaction energies summarized in table 1 show that both the recombinative CO2 adsorption ( eV) and abstraction ( eV) involving the O atom adsorbed atop Se (O hereafter) are highly exothermic. Recombination with the O atom adsorbed in the Se vacancy (O hereafter) is energetically less favored. In particular, adsorption is exothermic by only eV, while abstraction becomes endothermic, being the reaction energy eV. The different adsorption energy for O ( eV) and O ( eV) 51 is the main reason for the different reaction energies obtained with each adsorbate. CO adsorption near O although exothermic is energetically less favored than CO2 formation. In this adsorption configuration (denoted as CO in table 1), the CO center of mass is practically over a hollow site (see the black-framed panel in Figure 1). Interestingly, the calculated CO adsorption energy on 2O-MoSe2-vSe ( eV) is similar to that of CO adsorbed over a hollow site on pristine MoSe2 ( eV) 51. Nonetheless, since the focus of this work is the recombinative formation of CO2, we did not perform a systematic analysis of all possible CO adsorption configurations on 2O-MoSe2-vSe and, therefore, we cannot discard the existence of CO adsorption configurations more stable than CO.
Reaction | Reactant/Product Adsorption Site | (eV) |
CO + O CO | O/– | |
CO + O CO | O/– | |
CO + O CO | O/ CO | |
CO + O CO | O/ CO | |
CO CO | –/CO |
In addition to calculating the reaction energy, we carried out a preliminary inspection of the possible energy barriers the incoming CO may encounter while approaching the surface. For simplicity, we focused on exploring a few paths running close to O. In these calculations the height of the C atom (and the 3 coordinates of one Mo atom located more than 7 Å away the O adsorbates) are always fixed. In some paths, also the coordinates of the C and O atoms forming the molecule are fixed. The rest of degrees of freedom are allowed to relax until forces are smaller than 0.015 eV/Å. The dependence of the potential energy on along these inspection paths is represented in Figure 1. A vertically oriented CO approaching atop O will face an energy barrier of about 0.378 eV, if the CO molecule is forced to stay perpendicular to O by fixing the center of mass coordinates (green diamonds). The abrupt energy decrease at Å corresponds to abstraction of the adsorbed O to form a vertically oriented CO2 (see top and side views of the structures at points and in Figure 1). Note that the energy at is greater than that of desorbed CO2 (asymptotic limit of the CO2 curves plotted by blue and brown triangles in the figure), suggesting that an eventually formed vertical-CO2 will end desorbing. The latter was confirmed with an ulterior structural relaxation of the system configuration at , which ends with the CO2 molecule in vacuum. A similar energy barrier ( eV) is found when only the C coordinates (, , ) are fixed (red squares). In this case CO can tilt while approaching O (see structure in Figure 1) and also ends forming CO2 (see structure ), giving rise to the energy drop observed between points and in the figure. In this case the formed molecule is oriented parallel to the surface and it is energetically more stable than the desorbed CO2. Finally, when only is fixed while approaching over O, CO tilts and is progressively repealed away the initial O–C vertical line, avoiding the energy barriers found in the two previous paths. The minimum energy configuration along this path is shown in the black framed panel. For completeness, we also calculate potential energy curves against for the CO2 formed from CO+O. In these calculations only is fixed. Energies for CO2 approaching over O and atop the Se atom on which O was adsorbed prior recombining are respectively plotted by brown and blue curves in Figure 1. Corresponding minimum energy structures are depicted in the brown and blue framed panels.
Although CO2 formation and desorption upon O abstraction is energetically favored, the previous analysis casts some doubts regarding the likeliness of the reaction. The explored paths present energy barriers in the entrance channel that may difficult the required approach of the CO molecule to O prior recombining. Nevertheless, we cannot exclude the existence of other complex but more favorable reaction paths with lower or nonexistent energy barriers. In this respect, there are many examples showing that the incoming gaseous species can also capture the adsorbate upon reflecting once (indirect ER recombination) or several times (hot atom/molecule recombination) with the surface. Both direct and indirect ER recombination are observed in molecular dynamics (MD) and AIMD simulations of H scattered off Cl/Au(111) 68, 69 and H/Cu(111) 70, 71 and in MD simulations of N scattered off N-covered Ag(111) 72, 73 and N-covered W(110) 74, 75, 76, whereas only indirect ER occurs in the N-covered W(100) 74, 75 even at grazing incidence conditions at which direct ER is likely to occur 76 and, important for the present study, in the formation of CO2 by O atoms impinging on CO/Pt(111) 77. In the case of H impinging on H-covered W(110) and W(100) surfaces, all ER recombinations are indirect 76, being also important the hot atom mechanism as shown in ulterior simulations performed at finite coverage 78, 79, 80.
Considering the rich reactivity observed in other systems, the results of our AIMD calculations are quite surprising. Even if most of the simulated incidence energies ( eV) suffice to overcome the energy barriers explored in Figure 1, none of the trajectories led to CO2 formation. The latter includes both desorption and trapping (adsorption) of the expected CO+O recombination. There are only two possible outcomes observed at the end of simulations that correspond to the impinging CO being either reflected or molecularly trapped on the surface. Processes such as O atomic desorption or O2 recombinative desorption are highly endothermic and, therefore, energetically forbidden for the incidence energies here considered. Figure 2 shows the probabilities for CO reflection and trapping as a function of the incidence energy . At the lowest considered energy ( eV) 65% of the molecules are reflected, whereas a high 35% of them remain dynamically trapped at the end of the simulations (1 ps). Once the first collision with the surface takes place within a short interval of 2–4 fs, the trapped CO molecules experience a second rebound in the vacuum side at very diverse heights ( Å) and are attracted back to the surface. After extending the integration time up to about 1.4 ps, all these molecules stay below 6 Å from the surface, but the final outcome of these trajectories at longer times remains unclear with our limited-time dynamics (see Figure S1 in the Supporting Information). As increases, the trapping probability rapidly decreases in favor of CO reflection. Thus, at the largest eV more than 95% of the molecules are reflected. With very few exceptions, observed at lower incidence energies, the reflection process occurs after the very first collision with the surface. For eV, the first (classical) turning point of the reflected molecules is distributed in three regions depending on whether CO collides near O ( Å ), near O ( Å ), or at other points on the surface ( Å). These turning point regions move closer to the surface and broaden as increases. The turning point distributions for the trapped molecules exhibit a less featured structure. All these observations can be seen in Figures S2, S3, and S4 in the Supporting Information.
A more thorough analysis of the trajectories allows us to better understand the lack of recombination processes. Figures 3 and 4 show the evolution along the incoming trajectory of the CO center of mass coordinates () and the polar angle w.r.t. the surface normal of the CO molecular axis for incidence energies eV and eV, respectively. (Note in passing that the center of mass position at Å provides a clear idea of the area covered by the random distribution over the surface that was used in the initial conditions sampling described in sec AIMD Initial Conditions.) The number of reflected and trapped molecules that reach a certain height during the initial approach is written on each column, providing information on how many molecules have experienced the first collision with the surface between two consecutive height values. The most salient feature in both figures is the inability of the CO molecules to get close enough to O. In the case of eV, already at Å, molecules that were impinging nearby O have been reflected (6 molecules) or have been deviated to remain trapped at the surface (5 molecules). At the higher energy eV (Figure 4), the molecules can get closer. Yet two of the molecules incoming above O have already been reflected at 3.5 Å and at 3 Å all the molecules that were approaching O have been reflected or deviated from the O proximity in the case of the trapped molecules. Interestingly, a majority of the molecules that can get as close to the surface as Å are the ones that are in the region close to O. However, O is buried in the Se topmost layer ( Å) 51 and strongly bound to the substrate, such that recombination does not take place. Regarding the distributions, there is no noticeable effect related to the orientation of the molecules, except for a certain predominance of near parallel orientations () at the closest distances.
These results are consistent with the information obtained from the reaction paths analyzed in Figure 1. The reaction is highly exothermic and direct once the CO molecule is close to O, with the C atom at a distance from the surface below 3.5–4 Å depending on its orientation. However, the energy barriers that exist in the incoming path prevent the molecules from reaching that configuration. As a consequence, the CO molecules either reflect directly on top of O before getting close enough to react, or are deviated and remain trapped at the surface. Though we cannot completely exclude that the trapped molecules may eventually get close to the O in a proper orientation and react, the probability for this process is expected to be very low, considering that it has not occurred in any of the trapped trajectories for which the integration time has been increased up to 1.2–1.4 ps. In any case, in spite of its high exothermicity, direct ER recombination between CO impinging from vacuum and O can apparently be excluded. To further support this conclusion, one additional trajectory was run for each , in which the initial CO coordinates correspond to a vertically oriented CO located exactly above O at height Å. The turning point (taken as the height of the C atom) decreases to 2.49 Å for eV and to 2.03 Å for eV but in all cases the CO molecule bounces back into vacuum highly rotationally excited without forming a stable O–C bound.
Regarding the difference between reflected and trapped CO, Figures 3 and 4 suggest that those molecules initially impinging over or nearby O are more likely to be trapped after the first collision with the surface. But there is also a certain amount of trapped molecules that were initially located nearby O. The evolution of the molecular orientation features no apparent differences between reflected and trapped CO. Altogether the difference between direct reflection and trapping seems related to minor details of the potential energy felt by the incoming CO molecule.
Finally, focusing on the reflected trajectories, in the left panel of Figure 5 we observe that these molecules are predominantly reflected with small outgoing angles (). This is expected for molecules under normal incidence that are directly reflected, i.e., without experiencing any relevant deviation along the incoming trajectory. The latter suggests a negligible corrugation of the potential energy surface as compared to the incidence energy . In this respect, the wider distribution of the outgoing angle at eV agrees with the early energy barriers discussed above. Upon colliding with the surface, the reflected molecules become quite rotationally excited, particularly for incidence energies , and 1.0 eV. As shown in the right panel of Figure 5, the rotational quantum number reaches values as high as for the largest studied .
4 Conclusions
We have studied the interaction of CO molecules coming from vacuum and a 2O-MoSe2-vSe surface, in which one O atom is adsorbed in a Se vacancy and a second O atom is adsorbed on top of a Se atom of the phase of the MoSe2 monolayer. The latter is the equilibrium structure upon (dissociative) adsorption of a O2 molecule on a MoSe2 monolayer with one Se vacancy 51. Interestingly, DFT calculations show that the CO recombination with the O atom adsorbed on top of the Se atom with either subsequent adsorption or desorption of the formed CO2 molecules are highly exothermic reactions with energy gains of around 3–3.5 eV. Motivated by this finding, we have performed AIMD simulations with five different initial translational energies of CO molecules scattered off a 2O-MoSe2-vSe surface thermalized at 100 K. Our aim was to analyze how likely was the CO oxidation in this system and the corresponding reaction mechanisms. However, none of the 200 AIMD trajectories run for each energy has led to CO2 formation. The majority of the CO molecules are directly reflected, albeit in some cases (mainly at the lowest analyzed energies) some of them remain trapped at the surface without reacting with the adsorbed O atoms. The analysis of the trajectories show that those CO molecules impinging on the O are reflected or deviated to trapping before they can get close enough to the O atom in order recombination can occur. Inspection of the possible reaction paths suggests that the configurational space of the oxidation process via an Eley-Rideal process is very restrictive. Due to the unavoidable limited statistics and integration time, which does not allow to completely determine the final outcome of the trapped CO molecules, we can not absolutely exclude the possibility of the reaction to occur at a later stage. Nevertheless, we can safely conclude that, in spite of its large exothermicity, the recombinative CO+O abstraction is in the best case a very unlikely reaction in this system. As a general and important conclusion, the results of our dynamics simulations remark the importance of going beyond a pure static analysis of the reaction energetics to determine the catalytic character of any material.
Financial support provided by the Spanish MCINAEI10.13039501100011033, FEDER Una manera de hacer Europa (Grant No. PID2022-140163NB-I00), Gobierno Vasco-UPV/EHU (Project No. IT1569-22), and the Basque Government Education Departments’ IKUR program, also co-funded by the European NextGenerationEU action through the Spanish Plan de Recuperación, Transformación y Resiliencia (PRTR). We also acknowledge the HPC resources provided by the Donostia International Physics Center (DIPC) Supercomputing Center. R.B. acknowledges funding from ADAGIO (Advanced Manufacturing Research Fellowship Programme in the Basque – New Aquitaine Region) from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska Curie cofund Grant Agreement No. 101034379.
The Supporting Information is available free of charge at XXX and includes detailed information of the turning point for the reflected and trapped molecules.

### v2

@@
 […]
-![[Uncaptioned image]](extracted/5871042/figures/TOC-1.png)
 1 Introduction
@@
 Nonetheless, both theory and experiments reveal that the efficiency of these two reactions can be enhanced at the edges of the nanosheets 37, 36, 3, around chalcogen and metal vacancies 31, 32, 24, and by doping the nanosheet 33.
-Other chemical reactions that are the focus on theoretical studies of the TMD catalytic properties include, for example, CO2 reduction to produce acids and other valuable products 38, 39, 40, 41, 42, 43, 1, 2, 6, N2 reduction to produce ammonia 44, production of ethanol from syngas 45, and also CO oxidation on MoS2 doped with metal atoms 46, 47, on PtTe2 in the presence of O2 48 and on the Janus WSSe surface 49.
+Other chemical reactions that are the focus on theoretical studies of the TMD catalytic properties include, for example, CO2 reduction to produce acids and other valuable products 38, 39, 40, 41, 42, 43, 1, 2, 6, N2 reduction to produce ammonia 44, production of ethanol from syngas 45, and also CO oxidation on MoS2 doped with metal atoms 46, 47, on PtTe2 in the presence of O2 48, and on the Janus WSSe surface 49.
 In the aforereviewed bibliography, the catalytic properties are commonly determined in terms of the adsorption energies and minimum energy paths of the reaction considered.
-Being the latter the common procedure, the purpose and novelty of our work are to explore the catalytic properties of MoSe2 for CO2 formation but from a different perspective.
-In particular, we focus on the dynamics of the recombination process between one O atom adsorbed on the MoSe2 surface with a Se vacancy and a gas-phase CO molecule.
+Being the latter the common procedure, the purpose and novelty of our work are to explore the catalytic properties of MoSe2 for CO2 formation but from a different perspective, the reaction dynamics.
+DFT calculations including van der Waals corrections showed that the dissociative adsorption of O2, which is endothermic by more than 1 eV in the pristine 1H-MoSe2 sheet, becomes highly exothermic in the presence of Se vacancies 50.
+In light of those results, here we focus on the dynamics of the recombination process between one O atom adsorbed on the MoSe2 surface with a Se vacancy and a gas-phase CO molecule.
 This kind of reaction process, which can occur upon a single collision of the gas species with either the adsorbate (direct Eley-Rideal) or the surface (indirect Eley-Rideal) or after few collisions that serve to slow down the incoming atom/molecule (hot atom/molecule abstraction), are expected to efficiently compete with the thermal Langmuir-Hinshelwood recombination when the adsorbate is strongly bound to the surface.
 This is precisely the case in the dissociative adsorption of O2 on the MoSe2 surface with a Se vacancy.
-DFT calculations with the revPBE-vdW exchange-correlation functional developed by Dion et al.
-50 showed that the dissociative adsorption of O2 is exothermic, being the adsorption energies of the nascent O atoms that adsorb at different sites eV and eV 51.
+As shown in ref 50, the adsorption energies of the nascent O atoms formed upon the dissociative adsorption of O2 in that surface are as strong as eV and eV, depending on the adsorption site.
 Our new calculations for the recombinative formation of CO2 via an Eley-Rideal mechanism show that the reaction is energetically favored and thus, motivated us to study the process with ab initio molecular dynamics simulations (AIMD).
@@
 The paper is organized as follows, the general computational settings employed to both model the system and simulate the reaction dynamics with the AIMD methodology are detailed in the Theoretical Methods section.
-The Results and Discussion section starts with the analysis of the energetics of the main reactions that may occur in the system for the incidence conditions considered, i.e., a gas-phase CO impinging on the MoSe2 surface with a Se vacancy and an oxygen coverage of 0.125 ML.
-Next, the AIMD results are presented and discussed in the light of the information extracted from the in depth analysis of the trajectories.
+The Results & Discussion section starts with the analysis of the energetics of the main reactions that may occur in the system for the incidence conditions considered, i.e., a gas-phase CO impinging on the MoSe2 surface with a Se vacancy and an oxygen coverage of 0.125 ML.
+Next, the AIMD results are presented and discussed in light of the information extracted from the in depth analysis of the trajectories.
 The summary and conclusions of our work are given in the last section.
@@
 2.1 General DFT Computational Settings
-All DFT calculations are performed with the Vienna ab initio simulation package (vasp) 52 using the revPBE-vdW exchange-correlation functional proposed by Dion et al.
-50.
+All DFT calculations are performed with the Vienna ab initio simulation package (vasp) 51 using the revPBE-vdW exchange-correlation functional proposed by Dion et al.
+52.
 The ionic cores are described with the projector augmented wave (PAW) method 53 implemented in VASP 54.
@@
 Calculations are performed for the phase of the MoSe2 monolayer with a small concentration of non-interacting Se vacancies (denoted as MoSe2-vSe hereafter).
-Following ref 51, the latter is adequately described by a 44 surface supercell that contains a single Se vacancy and is separated from its periodic image by 23 Å of vacuum along the surface normal.
-The value of the hexagonal lattice constant is =3.386 Å 51.
-The O2 molecule is dissociatively adsorbed in the configuration analyzed in ref 51.
+Following ref 50, the latter is adequately described by a 44 surface supercell that contains a single Se vacancy and is separated from its periodic image by 23 Å of vacuum along the surface normal.
+The value of the hexagonal lattice constant is =3.386 Å 50.
+The O2 molecule is dissociatively adsorbed in the configuration analyzed in ref 50.
 Namely, one of the O atoms is adsorbed on top of the Se vacancy and the other one on top of the closest Se atom.
@@
 We denote this structure as 2O-MoSe2-vSe.
+Regarding our choice of the revPBE-vdW functional, our previous study of the adsorption properties of different diatomic molecules on MoSe2 and MoSe2-vSe showed that both revPBE-vdW and PBE-D3 56 provide a satisfactory description of the experimental MoSe2 geometrical and electronic structures.
+Furthermore, even if PBE-D3 tends to overbinding as compared to revPBE-vdW, there were no significant qualitative differences between both functionals in the O, O2, and CO adsorption and dissociation energies and adsorption sites on MoSe2 and MoSe2-vSe.
 2.2 AIMD Initial Conditions
@@
 The initial positions and velocities of the atoms in the substrate are randomly taken from a canonical ensemble of configurations.
-The latter is generated by thermalizing the 2O-MoSe2-vSe substrate at 100 K during more than 10 ps using the Nosé-Hoover thermostat 56 implemented in VASP.
-The use of constant-energy AIMD simulations with an appropriate set of initial positions and velocities constitutes a reasonable approximation to incorporate surface temperature effects during the short simulation time of interest (up to 1.0–1.4 ps in our case) 57.
-This is a common procedure in gas-surface dynamics simulations 58, 59, 60, 61, 57, 62, 63, 64, 65, 66, 67.
+The latter is generated by thermalizing the 2O-MoSe2-vSe substrate at 100 K during more than 10 ps using the Nosé-Hoover thermostat 57 implemented in VASP.
+The use of constant-energy AIMD simulations with an appropriate set of initial positions and velocities constitutes a reasonable approximation to incorporate surface temperature effects during the short simulation time of interest (up to 1.0–1.4 ps in our case) 58.
+This is a common procedure in gas-surface dynamics simulations 59, 60, 61, 62, 58, 63, 64, 65, 66, 67, 68.
 The CO molecule starts with its center of mass (CM) located at a distance Å from the 2O-MoSe2-vSe surface with the vibrational zero-point energy and zero rotational energy.
@@
 In the case of recombinative CO2 formation, the same criteria is used to distinguish between CO2 desorption and trapping.
-3 Results and Discussion
+3 Results & Discussion
 The calculated reaction energies summarized in table 1 show that both the recombinative CO2 adsorption ( eV) and abstraction ( eV) involving the O atom adsorbed atop Se (O hereafter) are highly exothermic.
@@
 In particular, adsorption is exothermic by only eV, while abstraction becomes endothermic, being the reaction energy eV.
-The different adsorption energy for O ( eV) and O ( eV) 51 is the main reason for the different reaction energies obtained with each adsorbate.
+The different adsorption energy for O ( eV) and O ( eV) 50 is the main reason for the different reaction energies obtained with each adsorbate.
 CO adsorption near O although exothermic is energetically less favored than CO2 formation.
 In this adsorption configuration (denoted as CO in table 1), the CO center of mass is practically over a hollow site (see the black-framed panel in Figure 1).
-Interestingly, the calculated CO adsorption energy on 2O-MoSe2-vSe ( eV) is similar to that of CO adsorbed over a hollow site on pristine MoSe2 ( eV) 51.
+Interestingly, the calculated CO adsorption energy on 2O-MoSe2-vSe ( eV) is similar to that of CO adsorbed over a hollow site on pristine MoSe2 ( eV) 50.
 Nonetheless, since the focus of this work is the recombinative formation of CO2, we did not perform a systematic analysis of all possible CO adsorption configurations on 2O-MoSe2-vSe and, therefore, we cannot discard the existence of CO adsorption configurations more stable than CO.
@@
 A vertically oriented CO approaching atop O will face an energy barrier of about 0.378 eV, if the CO molecule is forced to stay perpendicular to O by fixing the center of mass coordinates (green diamonds).
-The abrupt energy decrease at Å corresponds to abstraction of the adsorbed O to form a vertically oriented CO2 (see top and side views of the structures at points and in Figure 1).
-Note that the energy at is greater than that of desorbed CO2 (asymptotic limit of the CO2 curves plotted by blue and brown triangles in the figure), suggesting that an eventually formed vertical-CO2 will end desorbing.
-The latter was confirmed with an ulterior structural relaxation of the system configuration at , which ends with the CO2 molecule in vacuum.
+The abrupt energy decrease at Å corresponds to abstraction of the adsorbed O to form a vertically oriented CO2 (see top and side views of the structures at points a and b in Figure 1).
+Note that the energy at b is greater than that of desorbed CO2 (asymptotic limit of the CO2 curves plotted by blue and brown triangles in the figure), suggesting that an eventually formed vertical-CO2 will end desorbing.
+The latter was confirmed with an ulterior structural relaxation of the system configuration at b, which ends with the CO2 molecule in vacuum.
 A similar energy barrier ( eV) is found when only the C coordinates (, , ) are fixed (red squares).
-In this case CO can tilt while approaching O (see structure in Figure 1) and also ends forming CO2 (see structure ), giving rise to the energy drop observed between points and in the figure.
+In this case CO can tilt while approaching O (see structure c in Figure 1) and also ends forming CO2 (see structure d), giving rise to the energy drop observed between points c and d in the figure.
 In this case the formed molecule is oriented parallel to the surface and it is energetically more stable than the desorbed CO2.
@@
 In this respect, there are many examples showing that the incoming gaseous species can also capture the adsorbate upon reflecting once (indirect ER recombination) or several times (hot atom/molecule recombination) with the surface.
-Both direct and indirect ER recombination are observed in molecular dynamics (MD) and AIMD simulations of H scattered off Cl/Au(111) 68, 69 and H/Cu(111) 70, 71 and in MD simulations of N scattered off N-covered Ag(111) 72, 73 and N-covered W(110) 74, 75, 76, whereas only indirect ER occurs in the N-covered W(100) 74, 75 even at grazing incidence conditions at which direct ER is likely to occur 76 and, important for the present study, in the formation of CO2 by O atoms impinging on CO/Pt(111) 77.
-In the case of H impinging on H-covered W(110) and W(100) surfaces, all ER recombinations are indirect 76, being also important the hot atom mechanism as shown in ulterior simulations performed at finite coverage 78, 79, 80.
+Both direct and indirect ER recombination are observed in molecular dynamics (MD) and AIMD simulations of H scattered off Cl/Au(111) 69, 70 and H/Cu(111) 71, 72 and in MD simulations of N scattered off N-covered Ag(111) 73, 74 and N-covered W(110) 75, 76, 77, whereas only indirect ER occurs in the N-covered W(100) 75, 76 even at grazing incidence conditions at which direct ER is likely to occur 77 and, important for the present study, in the formation of CO2 by O atoms impinging on CO/Pt(111) 78.
+In the case of H impinging on H-covered W(110) and W(100) surfaces, all ER recombinations are indirect 77, being also important the hot atom mechanism as shown in ulterior simulations performed at finite coverage 79, 80, 81.
 Considering the rich reactivity observed in other systems, the results of our AIMD calculations are quite surprising.
@@
 At the lowest considered energy ( eV) 65% of the molecules are reflected, whereas a high 35% of them remain dynamically trapped at the end of the simulations (1 ps).
-Once the first collision with the surface takes place within a short interval of 2–4 fs, the trapped CO molecules experience a second rebound in the vacuum side at very diverse heights ( Å) and are attracted back to the surface.
+Once the first collision with the surface takes place within a short interval of 200–400 fs, the trapped CO molecules experience a second rebound in the vacuum side at very diverse heights ( Å) and are attracted back to the surface.
 After extending the integration time up to about 1.4 ps, all these molecules stay below 6 Å from the surface, but the final outcome of these trajectories at longer times remains unclear with our limited-time dynamics (see Figure S1 in the Supporting Information).
@@
 Interestingly, a majority of the molecules that can get as close to the surface as Å are the ones that are in the region close to O.
-However, O is buried in the Se topmost layer ( Å) 51 and strongly bound to the substrate, such that recombination does not take place.
+However, O is buried in the Se topmost layer ( Å) 50 and strongly bound to the substrate, such that recombination does not take place.
 Regarding the distributions, there is no noticeable effect related to the orientation of the molecules, except for a certain predominance of near parallel orientations () at the closest distances.
@@
 We have studied the interaction of CO molecules coming from vacuum and a 2O-MoSe2-vSe surface, in which one O atom is adsorbed in a Se vacancy and a second O atom is adsorbed on top of a Se atom of the phase of the MoSe2 monolayer.
-The latter is the equilibrium structure upon (dissociative) adsorption of a O2 molecule on a MoSe2 monolayer with one Se vacancy 51.
+The latter is the equilibrium structure upon (dissociative) adsorption of a O2 molecule on a MoSe2 monolayer with one Se vacancy 50.
 Interestingly, DFT calculations show that the CO recombination with the O atom adsorbed on top of the Se atom with either subsequent adsorption or desorption of the formed CO2 molecules are highly exothermic reactions with energy gains of around 3–3.5 eV.
@@
 The majority of the CO molecules are directly reflected, albeit in some cases (mainly at the lowest analyzed energies) some of them remain trapped at the surface without reacting with the adsorbed O atoms.
-The analysis of the trajectories show that those CO molecules impinging on the O are reflected or deviated to trapping before they can get close enough to the O atom in order recombination can occur.
+The analysis of the trajectories show that those CO molecules impinging on O are reflected or deviated to trapping before they can get close enough to the O atom in order recombination can occur.
 Inspection of the possible reaction paths suggests that the configurational space of the oxidation process via an Eley-Rideal process is very restrictive.